OpenStudy (anonymous):

Translate to a system of equations. Do not solve Two angels are supplementary. One angle is 3 degrees less than twice the other. Find the measures of the angles. Let x equal the first angle and y equal the second. x+y= x=

OpenStudy (anonymous):

x + y = 180 and x = 2y - 3

OpenStudy (anonymous):

Supplementary means 180 degrees..

OpenStudy (anonymous):

When you add the x and the y, you will get 180 because of the definition of supplementary angles.

OpenStudy (anonymous):

Now, you can solve by either substitution or elimination.

OpenStudy (anonymous):

Suppose x any y are the angles.. They are supplementary means, they add up to give 180 degrees.. \[x + y = 180^{\circ}\]

OpenStudy (anonymous):

I understand the whole 180 degrees what im not understanding is what x is

OpenStudy (anonymous):

And now, since you have for the second equation, an expression for x, I would suggest simply taking that second equation right-hand side and putting that into the first equation by substitution. "x" is 3 less than twice the other. Twice the other is 2y 3 less than that is 2y - 3

OpenStudy (anonymous):

So, x = 2y - 3

OpenStudy (anonymous):

Using x = 2y - 3 in the first equation x + y = 180 -> (2y - 3) + y = 180 You said do not solve, so that is the substitution, which can be solved for y because it is now one equation in one variable.

OpenStudy (anonymous):

And once you solve for "y", if you later do have to solve it, you take that "y" and put it back into the original equation #1 to get your "x".

OpenStudy (anonymous):

okay so what if the angle changed to say 5 degrees and four times less than four times the other

OpenStudy (anonymous):

I'm not sure I understand your last question the way it is written.

OpenStudy (anonymous):

sorry say if the degree of the angle changed to 5 degrees instead of 3 and its 5 degrees less than four times the other. does that make sense

OpenStudy (anonymous):

x has to be an equation.

OpenStudy (anonymous):

Then you would have a system of equations where the first equation is the same, but the second one changes: x + y = 180 and x = 4y - 5 The principle and problem resolution stays the same.

OpenStudy (anonymous):

The general form for the second equation here is: x = ny - m where we have "n" times the other angle and "m" less than that.

OpenStudy (anonymous):

Is that making sense to you now?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

You're welcome.